Which do you​ prefer: a bank account that pays 5.7% per year​ (EAR) for three years or

a. An account that pays 2.6 % every six months for three​ years?                   

b. An account that pays 7.2% every 18 months for three​ years?                 

c. An account that pays 0.28% per month for three​ years?

​(Note: Compare your current bank EAR with each of the three alternative accounts. Be careful not to round any intermediate steps less than six decimal​ places.)                                    

If you deposit $1 into a bank account that pays 5.7% per year for three​ years:The amount you will receive after three years is

​$
_____.

​(Round to five decimal​ places.)

a. An account that pays 2.6% every six months for 3​ years?                                         If you deposit $1

into a bank account that pays 2.6%

every six months for three​ years:The amount you will receive after three years is

​$______.

​(Round to five decimal​ places.)Which bank account would you​ prefer?

. ​(Select from the​ drop-down menu.)b. An account that pays 7.2%

every 18 months for 3​ years?                                          If you deposit $1

into a bank account that pay 7.2%

every 18 months for three​ years:The amount you will receive after three years is

$ .

​(Round to five decimal​ places.)Which bank account would you​ prefer?

c. An account that pays 0.28%

per month for three​ years?If you deposit $ 1

into a bank account that pays 0.28%

per month for three yearsThe amount you will receive after three years is

​$.

​ (Round to five decimal​ places.)Which bank account would you​ prefer?
Answer #1

A = P (1+ r/n) ^ nt

A = Amount, P = Principal, r = interest rate per annum

n= no. of times interest is compounded per year, t = No. of years

P = 1, r = 0.057, n = 1, t = 3

     = 1 ( 1+0.057)^1*3

      = 1.18093

If you deposit $1 into a bank account that pays 5.7% per year for three​ years:The amount you will receive after three years is

​$ 1.18093/-

P = 1, r = 0.026, n = 2, t = 3

= 1 (1+0.026)^(2*3)

= 1.16650

Answer to a : $ 1.16650/-              Would prefer earlier account with 5.7% p.a.

P = 1, r = 0.072, n = 12/18 = 0.67, t = 3

= 1 (1+0.072)^(0.67*3)

= 1.14998

Answer to b : $ 1.14998/-                  Would prefer this account as paying highest return.     

P = 1, r = 0.0028, n = 12 , t = 3  

= 1 (1+0.0028)^(12*3)

= 1.10590

Answer to c : $ 1.10590/-     Would prefer account with option b.